OpenStudy (timbo5lice):

A direct variation includes the point(7, 84)and (5, n) . Find n .

OpenStudy (timbo5lice):

@jordanjamesbay

OpenStudy (anonymous):

60

OpenStudy (anonymous):

7/84 = 5/n 7n = (84)(5) 7n = 420 n = 420/7 n = 60